train a is traveling at 100 mph and train b is traveling at 120mph. if train a passes a station at 10:25 and train b passes that station at 10:40. at what time will train b catch up to train a
Found 2 solutions by palanisamy, mido_007:
Answer by palanisamy(496)

Given, speed of the train a = 100mph
And, speed of the train b = 120mph
Given, train a passes a station at 10:25 and train b passes that station at 10:40.
Difference in time = 10:40-40:25 = 15minutes = 15/60 hour = 1/4 hour
Let the two trains meet after a time t hours after 10:40
Ditance travelled by b in t hours = 120t miles
Ditance travelled by a in (t+1/4) hours = 100(t+1/4) miles
When they meet, both distances are equal.
120t = 100(t+1/4)
120t = 100t+25
20t = 25
t = 25/20
t = 1 h 15m

Therefore the trains will meet at 10h40m+1h15m = 11h 55m

Answer by mido_007(1)

15/60*100=25 m
25/20= 1H 15m
time will be 10 40
+ 1 15
11 55
a will catch b at 11:55